OpenStudy (anonymous):

Can someone help me set this problem up i'm having a hard time In a right triangle, the hypotenuse is 1 inch longer than twice the shortest side. The third side is 7 inches longer than the shortest side. Find the length of each side.

jaynator495 (jaynator495):

Welcome To OpenStudy! Here you will find great helpers and friends, a community of students that help students! We hope you enjoy the experience!

jimthompson5910 (jim_thompson5910):

x = shortest side `the hypotenuse is 1 inch longer than twice the shortest side` so we know the hypotenuse is 2x+1 inches long `The third side is 7 inches longer than the shortest side` which means the third side is x+7 inches long you should have something like this drawing 

jimthompson5910 (jim_thompson5910):

agreed so far @Sooch ?

OpenStudy (anonymous):

@jim_thompson5910 Yes, I was getting the sides confused. so now i plug them into the a^2+b^2+c^2=0 formula?

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 actually

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

@jim_thompson5910 Got it! thank you so much

jimthompson5910 (jim_thompson5910):

no problem

jaynator495 (jaynator495):

Welcome To OpenStudy! Here you will find great helpers and friends, a community of students that help students! We hope you enjoy the experience!